### v1

@@ 4 Averaging Principle
Therefore, if $B$ is of polynomial growth, then so is $A$.
@@
5 Proof of the Main Theorem
With preparation above, we prove the main theorem now.
@@ 5 Proof of the Main Theorem
By Corollary 3.4 and [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
@@
Indeed, from (2.3),
Rearranging,
from which
@@
By integration-by-part formula,
$\mathbf{J}^{i}(\epsilon,t)$ (5.4)
$=$ $\epsilon\sum_{k}\int_{0}^{t}M^{-1}_{ik}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}[M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}dM^{-1}_{ik}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\epsilon\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}$
$=$ $\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$
@@
$\lesssim$ $\sum_{k}\mathbb{E}\Big(\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{2}ds\Big)^{p/2}$
$\lesssim$ $d\mathbb{E}\int_{0}^{t}|X^{\epsilon}_{s}-X_{s}|^{p}ds$
@@
($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
@@
We explore the limit of (5.14) as $\epsilon\to 0$ componentwise. For $\mathbf{A}(\epsilon,t),$ by (5.11), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of Proposition 3.1,
@@
where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ integrating by parts,
@@
$=$ $\epsilon\Big(\int_{0}^{t}X^{\epsilon}_{s}\otimes[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]\Big)_{ij}$
$=$ $\epsilon\int_{0}^{t}X^{\epsilon,i}_{s}[M^{-1}(X^{\epsilon}_{s})dY^{\epsilon}_{s}]^{j}$
$=$ $\epsilon\sum_{k}\int_{0}^{t}X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dY^{\epsilon,k}_{s}$
$=$ $\epsilon\sum_{k}\Big(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]\Big)$
$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t).$
By Proposition 3.1 and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
@@
On the other hand, notice that
$\mathbf{B}_{2}^{ij}(\epsilon,t)$ (5.18)
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}d[X^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})]$
$=$ $\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,i}_{s}+\epsilon\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}X^{\epsilon,i}_{s}\sum_{l}[\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})dX^{\epsilon,l}_{s}]$
$=$ $\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds+\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
$=:$ $\mathbf{B}_{21}^{ij}(\epsilon,t)+\mathbf{B}_{22}^{ij}(\epsilon,t).$
@@
$\mathbf{B}_{22}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.20)
@@
However, multipling both sides of (1.3) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right and then writing it componentwise, we see that
@@
which coincides with (5.21).

### v2

@@ 4 Averaging Principle
Therefore, AA has polynomial growth whenever $B$ does.
@@
5 Proof of Main Theorems
With preparation above, we prove the main theorems now.
@@ 5 Proof of Main Theorems
By Corollary 3.3 and interpolation theorem [9, Theorem A.15], it suffices to prove the pointwise convergence, that is,
@@
From (3.3)
@@
From (3.8) we see that
$\mathbf{J}^{i}(\epsilon,t)=\epsilon\sum_{k}M^{-1}_{ik}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}\partial_{x_{l}}M^{-1}_{ik}(X^{\epsilon}_{s})ds.$ (5.3)
@@
Assumptions ($\mathbf{A3}$) and ($\mathbf{A4}$) imply that
@@
We explore the limit of (5.13) as $\epsilon\to 0$ componentwise. For $\mathbf{A}(\epsilon,t),$ by (5.10), Burkholder–Davis–Gundy inequality and dominated convergence theorem with the help of (3.23),
@@
where the limit is in $L^{p}(\Omega,\mathcal{F},P).$ For $\mathbf{B}(\epsilon,t),$ we see from (3.17) that
@@
$=$ $\epsilon\sum_{k}X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t}-\sum_{k}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,i}_{s}M^{-1}_{jk}(X^{\epsilon}_{s})ds-\sum_{k,l}\int_{0}^{t}Y^{\epsilon,k}_{s}Y^{\epsilon,l}_{s}X^{\epsilon,i}_{s}\partial_{x_{l}}M^{-1}_{jk}(X^{\epsilon}_{s})ds$
$=:$ $\mathbf{B}^{ij}_{1}(\epsilon,t)-\mathbf{B}^{ij}_{2}(\epsilon,t)-\mathbf{B}^{ij}_{3}(\epsilon,t).$
By Proposition 3.1, (3.23) and ($\mathbf{A3}$), $(X^{\epsilon,i}_{t}M^{-1}_{jk}(X^{\epsilon}_{t})Y^{\epsilon,k}_{t})_{0<\epsilon\leq 1}$ is bounded in $L^{p}(\Omega,\mathcal{F},P)$ for each $t\in[0,T],$ so that
@@
$\mathbf{B}_{3}^{ij}(\epsilon,t)\to\sum_{k,l}\int_{0}^{t}X^{i}_{s}(\partial_{x_{l}}M^{-1}_{jk})(X_{s})J_{kl}(X_{s})ds,$ (5.18)
@@
However, multipling both sides of (1.9) by $M^{-1}$ on the left and $(M^{-1})^{\top}$ on the right handside and then writing it componentwise, we see that
@@
which coincides with (5.19).∎
Before demonstrating Theorem 2.4, we discuss a slightly more general result about transferring RDE to SDE. Consider the following RDE
$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{Y}_{t},$
where $\mathbf{Y}=(Y,\mathbb{Y})$ is a $\mathscr{C}^{\alpha}([0,T];\mathbb{R}^{d})$-valued random element, $Y$ is a continuous semimartingale,
[…]
the later integral being Itô integral, and $a$ is a $C([0,T];\mathbb{R}^{d\times d})$-valued random element. Then for almost every sample path, we have
$Z_{t}-Z_{s}$ (5.21)
$=$ $b(Z_{s})(t-s)+\sigma(Z_{s})Y_{s,t}+D\sigma(Z_{s})\sigma(Z_{s})\mathbb{Y}_{s,t}+R_{s,t},$
with $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|R_{s,t}|=0,$ $P$-a.s. . But
$(D\sigma(Z_{s})\sigma(Z_{s})\mathbb{Y}_{s,t})_{i}$ (5.22)
$=$ $\sum_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{s})\sigma_{kl}(Z_{s})\mathbb{Y}^{lj}_{s,t}$
[…]
where $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|\hat{R}_{s,t}|=0,$ $P$-a.s. . Combining (5.21) and (5.22),
$(Z_{t}-Z_{s})_{i}$
$=$ $[b_{i}(Z_{s})+\sum_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{s})\sigma_{kl}(Z_{s})a^{lj}_{s}](t-s)$
[…]
where $\lim\limits_{|\mathcal{P}|\to 0}\sup\limits_{\mathcal{P}\subset[0,T]}\sum\limits_{[s,t]\in\mathcal{P}}|\tilde{R}^{i}_{s,t}|=0,$ $P$-a.s. . By equivalence theorem of RDE and SDE[10, Proposition 6.9], for each $1\leq i\leq d,$
$dZ^{i}_{t}=[b_{i}(Z_{t})+\sum\limits_{j,k,l}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})a^{lj}_{t}]dt+\sum_{j}\sigma_{ij}(Z_{t})dY^{j}_{t}.$ (5.23)
Proof of Theorem 2.4:
By Theorem 2.2 and a probabilistic version of continuity of Itô–Lyons map[10, Theorem 6.1], $Z^{\epsilon}\to Z$ in $C^{\alpha}([0,T];\mathbb{R}^{d})$ in probability, with
$dZ_{t}=b(Z_{t})dt+\sigma(Z_{t})d\mathbf{X}_{t}\,.$
Now we transfer this RDE to SDE. By the same argument leading to (5.20), we have
$\mathbb{X}^{ij}_{s,t}$ (5.24)
$=$ $\Big(\int_{s}^{t}X_{u}\otimes\circ dX_{u}\Big)_{ij}+\frac{1}{2}\Big(\int_{s}^{t}J(X_{u})(M^{-1})^{\top}(X_{u})-M^{-1}(X_{u})J(X_{u})du\Big)_{ij}$
$=$ $\sum_{k,l}\int_{s}^{t}X^{i}_{u}S_{j}(X_{u})du+\sum_{k}\int_{s}^{t}X^{i}_{u}M^{-1}_{jk}(X_{u})F_{k}(X_{u})du$
$+\sum_{k}\int_{s}^{t}X^{i}_{u}M^{-1}_{jk}(X_{u})dW^{k}_{u}+\sum_{k}\int_{s}^{t}J_{ik}(X_{u})M^{-1}_{jk}(X_{u})du$
[…]
Applying (5.23),
$dZ^{i}_{t}$
$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})]dt+\sum_{j}\sigma_{ij}(Z_{t})dY^{j}_{t}$
$=$ $[b_{i}(Z_{t})+\sum\limits_{j,k,l,k^{\prime}}\partial_{z_{k}}\sigma_{ij}(Z_{t})\sigma_{kl}(Z_{t})J_{lk^{\prime}}(X_{t})M^{-1}_{jk^{\prime}}(X_{t})]dt$
$+[\sum_{j}\sigma_{ij}(Z_{t})S_{j}(X_{t})+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})F_{k}(X_{t})]dt+\sum_{j,k}\sigma_{ij}(Z_{t})M^{-1}_{jk}(X_{t})dW^{k}_{t}.$